i need help solving this problem:
(x+3.25)2=(x-4.1)2
this is what i have so far:
(x+3.25)(x+3.25)=(x-4.1)(x-4.1)

Answer by rfer(16322)

x^2+3.25x+3.25x+10.56=x^2-4.1x-4.1x+16.81
x^2+6.5x+10.56=x^2-8.2x+16.81
The x^2's cancel.
14.7x=6.25
x=6.25/14.7
x=0.425